Three girls shared a sum of money. 1st girl received ¼ of the amount, 2nd girl received ⅜ of the amount and the 3rd girl received $20 of the amount. what is the amount shared?
Answer by Boreal(15235)

a=1/4 x
b=3/8x
c=3/8 total=20

That means total is 20/(3/8)=160/3=$53.33 ANSWER
girl B also got $20.
girl A got $13.33, which is 1/4 of $53.33 to rounding.